OpenStudy (anonymous):

Prove:(tan x + sec x)^2 = 2sec^2 x + 2tan x sec x - 1 I tried it like 3 times

OpenStudy (anonymous):

6x

OpenStudy (anonymous):

@kirak95197 I am positive that is not correct

OpenStudy (anonymous):

Well your wrong

OpenStudy (anonymous):

F­uck you

OpenStudy (anonymous):

(secx + tanx)^2 = sec^2x + 2 secxtanx + tan^2x From the right side: 2sec^2x - 1 + 2secxtanx = sec^2x + (sec^2x - 1) = sec^2x + tan^2x + 2 secxtanx Thus it's proved!